a. Identify the limiting reagent when 15.0 g of H2 reacts with 48.0 g of O2 to produce water.
b. What is the mass of water produced?
1
Expert's answer
2013-01-15T08:24:39-0500

2H2 + O2 = 2H2O

n(H2) = m/M

n(H2) = 15g/2g/mol = 7,5 mol

n(O2) = 48g/32g/mol = 1,5 mol

a. The limiting reagent is O2

n(H2O) =1,5 mol

M(H2O) = 2*1 + 16 = 18 g/mol

m(H2O) = 18g*1,5 mol = 27g

b.The mass of water produced is 27g.